OpenStudy (anonymous):

9 3/4 percent coverted to fraction equals : ?????

OpenStudy (anonymous):

\[9 \frac{3}{4}=\]\[9+\frac{3}{4}=\]\[\frac{9\times4}{4}+\frac{3}{4}=\]\[\frac{36}{4}+\frac{3}{4}=\]\[\frac{36+3}{4}=\]\[\frac{39}{4}\]

OpenStudy (anonymous):

\[ \frac{39}{4}\]

OpenStudy (anonymous):

thanks